OpenStudy (anonymous):

A car travels 416 miles in 8 hours. At that rate, how far will the car travel in 22 hours? A.104 Miles B. 151.3 miles C. 1,144 miles D. 3,328 miles

OpenStudy (amistre64):

how many 8s are in 22?

OpenStudy (anonymous):

2

OpenStudy (amistre64):

2 and a smidge ... lets say 3 400 for every 8, so 3*400 = 1200 or a little less

OpenStudy (anonymous):

what is d? what is r? what is t d=rt

OpenStudy (anonymous):

d= 416 t = 8 r = d/t then, if r = _____, what is d at 22 hours?